Find the term of the geometric square.
The first term of a geometric sequence is 5 and the common ration is 3. Write the explicit formula for the sequence. Whats the 12th term of the sequence.

thank you for your help need answers ASAP i appreciate it

Answer by richwmiller(17219)

One question at a time.
Submit in proper categories.
Everyone wants everything ASAP or sooner.
What makes you special?
We know you waited till the last minute. What else?
Tn = 5 * 3^(n - 1)
T12 = 5 * 3^(11)
T12=5*177147=885735